### v1

# Words and Action: Modeling Linguistic Leadership in #BlackLivesMatter Communities

Abstract
In this project we describe a method of modeling semantic leadership across a set of communities associated with the #BlackLivesMatter movement, which has been informed by qualitative research on the structure of social media and Black Twitter in particular. We describe our bespoke approaches to time-binning, community clustering, and connecting communities over time, as well as our adaptation of state-of-the-art approaches to semantic change detection and semantic leadership induction. We find substantial evidence of the leadership role of BLM activists and progressives, as well as Black celebrities. We also find evidence of the sustained engagement of the conservative community with this discourse, suggesting an alternative explanation for how we arrived at the present moment, in which “anti-woke” and “anti-CRT” bills are being enacted nation-wide.
Introduction
Battles over social and political change are always waged with words and action–one need only look to the “Stop W.O.K.E. Act,” recently signed into Florida law, which prohibits the discussion of race and racism in public school and university classrooms, for evidence of how words index larger concepts and debates (?). While this particular bill was designed to censor discourse, ample evidence exists of how words can open up conversations, introduce new ideas into public consciousness, and impact social behavior and government policy (??). In this paper, we turn to one of the most visible contemporary examples of the intersection of words and action–the Black Lives Matter movement–to explore how words enter into a specific sociopolitical conversation, how they transform as that conversation evolves, and who is responsible for carrying forward those new or changed meanings.
Figure 1: Semantic leadership in #BlackLivesMatter Nodes represent the twelve main communities of the BLM network. The thickness of the edges are proportional to the number of word changes shared by each pair.
While there exists abundant research on Black Lives Matter (BLM), both as a hashtag (?????) and a movement (???), there are still open questions about how precisely BLM was able to “push the mainstream public sphere on issues of social progress” (?, p. xxvii). This is largely due to the unique community structure of Twitter, in which communities are defined not by individual users but by hashtags, and by the conversations among users that hashtags enable (????). Operationalizing these medium-specific theories about community formation on Twitter, as well as about “Twitter time,” we present a multipart model that allows us to 1) detect the latent communities associated with the #BlackLivesMatter hashtag and link them over time; and 2) detect semantic changes associated with each community using word embeddings. Together with a measure of semantic leadership (?), we are able to induce a semantic leadership network (see Figure 1), allowing us to identify the communities responsible for introducing new or changed word meanings into the BLM network and the communities responsible for adopting (or co-opting) them.
We find that the community of BLM activists plays an outsized role in introducing semantic changes into the network, and that these changes are most consistently taken up by the conservative community. We find a similar albeit weaker signal between the progressive and conservative communities, confirming the role of left-leaning discourse in shaping the general terms of debate (?). We also find that Black celebrities significantly shape the discourse, introducing word changes that the center/left news media as well as the conservative community later take up. The scale of our data, combined with our methods of validation, provide a new layer of evidence to support the largely qualitative scholarship that affirms the role of the BLM movement for bringing ideas about racial justice from the margin to the mainstream. Our results also provide early evidence of the conservative backlash that is now being expressed in a range of “anti-woke” legislation and related policies and political groups, since we find that the conservative community is the largest follower of all word changes — not only after the summer of 2020 but from the movement’s very beginning.
To summarize, our contributions are as follows:
  • Evidence at scale of how BLM activists shaped the discourse around racial justice, pushing new ideas from the margin to the mainstream
  • Early examples of specific terms (e.g. theory) that have since become central to government policy and debate
  • A precise network structure of the communities that comprise the #BlackLivesMatter hashtag, derived from extensive hand-labeling and domain-specific research
  • An example of how domain-specific theory–here, media studies scholarship on Black Twitter–can inform technical modeling approaches
Background and Prior Work
Research on Black Twitter
The topic of Twitter, and Black Twitter in particular, has been the subject of significant research in the fields of communication and media studies, among other humanities and social science fields. In Distributed Blackness (?), André Brock theorizes the community structure of Black Twitter as consisting of the people who are online and in conversation with each other at any given time. He also theorizes “Twitter time” as non-linear, defined more by specific events and the bursts of conversation they prompt than by any standard timescale. In Hashtag Activism (?), Sarah J. Jackson, Moya Bailey, and Brooke Foucault Welles propose a similar conception of Twitter communities organized by the hashtag itself. Hashtags, they explain, ”designate collective thoughts, ideas, arguments, and experiences that might otherwise stand alone.” Additional work on Black Twitter has considered the perspectives of the users themselves (?), its historical antecedents (?), and its possible futures (?). This qualitative work constitutes the foundation for our project. Our contribution represents an attempt to formalize the concepts and structures these scholars describe.
Online Social Movements
Several characteristics of online social movements, including the motivation and role of participants (???), activist influence on potential recruits (?), formation and sustenance of collective identities (?),and mobilization of activists (?????), have been extensively studied. BLM, in particular, has been studied as a unique example of a highly contested social movement with a sizeable online footprint (????). The online discourse of this movement has been shaped by tragic offline events (?) and has persisted for almost a decade (?). In terms of our project, the closest related work is Freelon, McIlwain, and Clark’s ”Beyond the Hashtags” report (?), which employs community detection followed by hand-labeling by domain experts in order to trace the changes in network structure over the first 18 months of the BLM movement. Recognizing the expertise involved in that project, our goal was to carry forward the authors’ work. Here, we extend the timeframe of analysis by another five years; develop new methods that allow us to expand the initial set of communities labeled by that team; and introduce an additional layer of linguistic analysis that complements the original study.
Linguistic methods have also been used to study different aspects of BLM. For example, the dynamics of participation have been shown to be correlated with textual features of tweets (?); the impact of police killings has been shown to reflect in the emotional narratives of social media participants (?); hashtag use has been shown to draw attention (??) or push back against issues (?). Framing and rhetorical strategies have been shown to be means of differentiation and coalition forming (???). We extend this line of work by inducing the leadership structure within the coalition of communities through the adoption of linguistic changes by community members.
Table 1: Descriptive statistics User and tweet statistics for each temporal partition are provided. Start and end dates are inclusive for each period.
Period Time Span Days Pre-filtered Tweets Post-filtered Tweets Post-filtered Users
1 Jan 01, 2014 - Sep 01, 2014 244 7,852,863 1,809,274 331,479
2 Sep 02, 2014 - Dec 24, 2014 114 11,590,864 2,744,459 497,168
3 Dec 25, 2014 - Aug 30, 2015 249 6,775,714 1,804,867 228,935
4 Aug 31, 2015 - Jul 07, 2016 312 2,539,907 695,180 143,220
5 Jul 08, 2016 - May 26, 2020 1419 7,196,284 1,630,376 347,567
6 May 27, 2020 - Aug 06, 2020 72 41,593,336 5,252,117 1,659,510
7 Aug 07, 2020 - Dec 31, 2020 146 5,112,396 877,086 214,050
Semantic Change and Leadership
Sociocultural changes often get encoded as lexical semantic changes in language (?). Consequently, computational methods to automatically extract semantic changes, given a timestamped text collection, have been developed with increasing regularity (??????, interalia). Semantic changes detected by such methods are demonstrably effective in two ways: 1) as first-order objects of analytical interest, such as in tracking the evolution of political issues (?) or shift in racial attitudes over time (?); 2) as units that can be aggregated to recover the latent leadership structure among scholarly articles (??) or newspapers (?). In this work, we follow the latter approach, using word embeddings to unveil the leadership dynamics between communities and their role in shaping online discourse of the BLM movement.
Figure 2: Volume of tweets per month We visualize the temporal distribution of tweets in our dataset in the style of Freelon et al (?). The seven temporal partitions are indicated here both by color and a reference label (e.g., Period 1). Some pivotal events to the BLM movement in each partition are annotated.
Data
Our dataset consists of tweets containing ”#BlackLivesMatter” between June 1, 2014 and December 31, 2020. These tweets were collected in two phases: tweets between June 1, 2014 and May 31, 2015 were rehydrated in April 2021 using twarc [^1] https://twarc-project.readthedocs.io/en/latest/ from tweet IDs provided by Freelon et al. ( ?). [^2] In addition to the #BlackLivesMatter hashtag, these tweets also included 44 additional ”keywords related to BLM and police killings of Black people under questionable circumstances” (?, p. 21) Our choice of this data source was motivated by our desire to build upon the hand-labeled communities in that paper, which we then used to seed our own community detection step (see next section). Tweets between June 1, 2015 and December 31, 2020, were collected using Twitter’s search API between April and August 2021. Together, this resulted in a total of 82,661,364 tweets from 14,914,359 unique users.
Filtering
Our analysis proceeds in two steps: a community detection step (see Community Detection and Matching) and a semantic leadership network induction step (see Leadership Network Induction). For the former, we filtered out all tweets tagged as non-English using langid (?); for the latter, we did additional filtering passes to remove tweets that were retweets, exact duplicates, or tagged as non-English by langid (?) and fasttext (?). The first round of filtering left 69,560,572 tweets; the second round of filtering left 32,652,123 tweets. The breakdown of tweets by timebin and unique users is shown in Table  1.
Ethics and Privacy
We remain attentive to concerns about extracting knowledge and intellectual labor from online communities for academic research (?????) throughout our analysis. These concerns motivate our modeling approach, which is designed to extend rather than replace existing scholarship. In addition, our multiracial project team has members who have engaged with the BLM movement in various roles, on and offline. With respect to concerns about privacy and attribution, we are not required to identify individual users in most cases, as the majority of our analysis is undertaken at an aggregate level. In the four instances where we identify individual users or their tweets, we take a context-specific approach (?). 6 requires that we give a sense of the users that comprise each community. To avoid unwanted exposure, we rank users by in-degree and name only those within the top five of each community that a) still have public accounts on X; and b) meet the ”reasonably public” threshold of 3000 or more followers as formulated by Freelon et al. (?). Table 2 requires that we provide examples of the word changes that our model detects. Here, we followed current best practice by paraphrasing tweets rather than quoting directly (??) because we do not want to expose activist users or users with low follower counts to unwanted attention or possible harm (?). We also confirm via X that our ”ethically fabricated” language (?) is not traceable back to the original user. The two users mentioned by name in the paper are public figures and as such, have upwards of one million followers.
Community Detection and Matching
Our first goal is to find relatively persistent communities of Twitter users throughout the full timespan of this study. In this section we describe the procedure for partitioning the timespan, mapping nodes to clusters within temporal partitions, and matching clusters across temporal partitions to find cohesive communities.
Defining Discrete Time Periods
We partition the dataset around the local maxima of tweet frequency, allowing irregular intervals. We find that the relative maxima correspond to real-world events relating to BLM. For instance, the first local maximum is found on August 11th, 2014, two days after the murder of Michael Brown  (??). We used these real-world events as a qualitative baseline for validating possible maxima. When additional maxima occur within 30 days of another maximum, they are grouped into one period (see Table 3). From the 11 relative maxima that were extracted using this method, six became markers for time periods resulting in seven intervals (see Figure 2 and Table 1).
Time-Specific User Clustering
User graphs
We use Freelon’s Twitter Subgraph Manipulator (TSM) [^3] https://github.com/dfreelon/TSM package to construct graphs from three forms of user-to-user interaction on Twitter found in tweets: replies, mentions, and retweets. These graphs are used to partition users into disjoint clusters in each of the seven temporal bins.
Clustering objective
To obtain user clusters, we maximize intra-group modularity (??) by modifying the Louvain clustering algorithm (?) — a well-known algorithm for detection of an unconstrained number of communities — using the TSM package. Because of our goal of extending the communities identified in Freelon, McIlwain, and Clark (?), which were hand-labeled by experts on BLM and Black Twitter, we augment the basic Louvain approach with six seed communities, consisting of 60 users, which were reconstructed from the communities documented in the original study. This improved the overall coherence of the clusters, which initial experiments had demonstrated to be poor.
Clusters
We extract the top 50 clusters for each temporal bin, which can be seen as slices of the larger BLM network, in the sense that they capture the time-bound representation of the BLM community structure.
Figure 3: Temporal tweet distribution The timeseries of tweet volume per community after filtering.
Figure 4: Community size Count of unique users in each community. Bar segments represent subcommunity clusters with height proportional to cluster size.
From Clusters to Communities
To find persistent communities from the ephemeral clusters obtained in the previous step, we cast the problem as a matching problem in a bipartite graph. Specifically, for a pair of successive intervals, we construct a bipartite graph linking each cluster from one interval to every cluster in the next interval. The weight on these links represents similarity between the two clusters, and is calculated as a weighted Jaccard score taking into account the in-degree of the users in a cluster (?). Following Freelon et al., we experimented with pruning the “weaker” edges if their similarity was below 0.3. However, since our intervals span over months or years, we see a greater level of fragmentation at this threshold. Consequently, we used a knee plot, which highlights the tradeoff between the number of matched communities and the jaccard threshold. Using this diagnostic, we lower the threshold to 0.07, a brightline that permits persistent communities with noise confined to spurious matches (see Figure 7).
We repeat the matching process between each successive temporal interval, and throw out any intermediary edges with similarity less than 0.07, effectively identifying the chains of clusters that remain most persistent over time. In summary, the sets of disjoint, independent clusters are matched to the preceding and succeeding periods producing 65 cohesive and persistent communities (see Figure 4).
Labeling
Labeling Users
With the sequenced slices matched, we can begin to develop qualitative characteristics for the 65 communities that result. We sought to generate a basic profile for each by surveying the top 20 users ranked by in-degree in each community, and giving each label from a set of iteratively redefined labels. The decision to use the top 20 users was heuristic; on average, the top 20 users made up 46.7% of the total in-degree, demonstrating their centrality to each of the 65 communities. A sub-label is included which qualifies the sort of membership a user has. For instance, a CNN journalist like @donlemon would be classified as “Established Media”, sub-labeled “Journalist.” Each user is given a second optional label if the first is insufficiently representative; these labels are not hierarchical, as they are jointly important. From this set of classifications and from background research, we agree upon a set of categories which could maximally encompass our communities.
In order to establish label agreement, the annotators constructed a series of pilot tests. These pilot test annotations yielded 83% agreement between annotators on average. With this established, each of the four annotators labeled 360 users. We randomly selected six communities (120 users) to be annotated by two annotators for agreement. Thus, annotators labeled 1,200 total users once and 120 twice to check for agreement. We computed Fleiss’s Kappa from the random pairwise assignment. Since some users received two non-ordinal labels, we present the maximum agreement score as reference. On average, interrater agreement was 0.798. After coming to consensus on how to define the mismatched labels, we continued from this user-level analysis to consolidating labels for the communities more broadly.
Describing and Grouping Communities
To develop a stronger sense of each community as a whole, the annotators used the set of labels assigned to its top 20 users to define a short description of each community. Using the two to four alternative descriptions for each community, the annotators then agreed on a final short description. In order to conduct semantic change analysis and perform permutation tests, it is necessary to condense the number of communities analyzed. Using the final short descriptions, we merged our 65 communities into 12 larger groups (see Figure  4 and Figure  6; ”miscellaneous” not shown).
We excluded two groups from this analysis, one for not being in English and the other for containing spam. Since communities on Twitter change in membership over time, it is important to check if these group labels are apt. Once grouped into the 12 macro communities, the average in-degree of the top 20 users was no less than 20%, suggesting that, relative to the size of their respective communities, these users play a central role in its dialogue. Thus, their presence provides important information about the background of other users and potential insight into how they may approach discussing social issues.
Leadership Network Induction
We adopted the method from ? (?) to induce a semantic leadership network between document sources. Soni et al. considered a set of historical newspapers as document sources and individual articles as documents. In contrast, we consider our Twitter user communities as document sources and individual tweets as documents.
Besides this superficial difference in the setup, there are two substantive differences. First, Soni et al. divided their timestamped collection of newspaper articles into temporal slices of equal timespans. In contrast, our tweet collection is divided into intervals demarcated by timestamps of real-world incidents. Second, Soni et al. considered document sources to remain fixed over time. In contrast, we consider communities of Twitter users as document sources which may undergo changes as a result of users joining, leaving and migrating between communities. We thus apply the methodology to our data but note that the interpretation of influence across communities should be qualified to account for these user dynamics.
For completeness, we briefly describe the core methodology here. Consider any tweet $i$ as a sequence of tokens (words) $w_{i}=(w_{i,1},w_{i,2},\ldots,w_{i,n_{i}})$ from a finite vocabulary $\mathcal{V}$, where $n_{i}$ indicates the tweet length. Our corpus consists of $N$ tweets, $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{N}\}$. Each tweet has two labels: a discrete timestamp $t_{i}$, obtained by binning the document timestamps into one of $T$ bins; and a community label $s_{i}$ based on the community of the tweet’s author.
Semantic Change Detection
To identify semantic changes, we learn temporal word-type embeddings from data. Other methods to detect semantic changes using token embeddings have been proposed but with mixed success (?). Consequently, we approach the task as learning word embeddings and then enhancing them to account for information from temporal or other facets. Words are mapped to embeddings, typically by maximizing the following probability under the skipgram language model (?) between any pair of nearby tokens $(w_{j},w_{j^{\prime}})$ in a single document ,
$P(w_{j^{\prime}}\mid w_{j})\propto\exp\left(\mathbf{v}_{w_{j^{\prime}}}\cdot\mathbf{u}_{w_{j}}\right),$ (1)
where $\mathbf{v}_{w_{j^{\prime}}}$ is the “output” embedding of $w_{j^{\prime}}$, and $\mathbf{u}_{w_{j}}$ is the “input” embedding of $w_{j}$; both set of embeddings are the parameters of the skipgram model that are learned from data. If documents are timestamped, then every input embedding can be split into a core embedding and a time-specific deviation as follows,
$\mathbf{u}_{w_{i,j}}^{(t_{i})}=\mathbf{b}_{w_{i,j}}+\mathbf{r}_{w_{i,j}}^{(t_{i})},$ (2)
where $\mathbf{b}_{w_{i,j}}$ is the base embedding of the word $w_{i,j}$ and $\mathbf{r}_{w_{i,j}}^{(t_{i})}$ is the residual for time $t_{i}$. The residual turns the core meaning to a temporally specific meaning and are hence regularized towards zero to follow the assumption that for most words the core meaning remains intact over time. Embedding decomposition of this type has been used in other applications such as to identify geographic variation (?) or perceptual differences in meaning (?).
Once the base and the residual components of the input embeddings for each word are obtained, semantic changes can be found by comparing near-neighbors in the embedding space (?). The upshot, at the end of this step, is a ranked order list of triples, consisting of the word that has changed, the timestamp for the onset of the change, and the timestamp for the conclusion of the change.
Identifying Semantic Leaders
For each semantic change, we then score the communities that lead or lag that change. To do this, the input embeddings in Equation 3 are modified to include a residual term for the community of each token. The input embedding is rewritten,
$\mathbf{u}_{w_{i,j}}^{(t_{i},s_{i})}=\mathbf{b}_{w_{i,j}}+\mathbf{r}_{w_{i,j}}^{(t_{i})}+\mathbf{r}_{w_{i,j}}^{(t_{i},s_{i})},$ (3)
where $\mathbf{r}_{w_{i,j}}^{(t_{i},s_{i})}$ is the source-specific temporal deviation added to the temporal and atemporal components of the input embedding.
Next, a leadership score is calculated between a pair of communities $s_{1}$ and $s_{2}$ and for a given change $(w,t_{1},t_{2})$, where $t_{1}<t_{2}$ are the timestamps of a change in the meaning of word $w$, as a ratio of cross-correlation between the sources to the auto-correlation in meaning as follows,
$\textsc{Lead}(s_{1}\rightarrow s_{2},w,t_{1},t_{2})=\frac{\mathbf{u}_{w}^{(t_{1},s_{1})}\cdot\mathbf{u}_{w}^{(t_{2},s_{2})}}{\mathbf{u}_{w}^{(t_{1},s_{2})}\cdot\mathbf{u}_{w}^{(t_{2},s_{2})}}$ (4)
The tuple $(s_{1}\rightarrow s_{2},w,t_{1},t_{2})$ can be referred as a leadership event denoted by $e$. A higher $\textsc{Lead}(e)$ score indicates more leadership; a score of $1$ corresponds to a baseline case of no leadership.
Inducing Leadership Network
We can aggregate the leadership scores for each word obtained from Equation 4 to construct a dense network between the communities. However, such a network may include changes that are, at best, spurious correlations due to random noise or structural biases such as temporal precedence of certain communities over others. To account for spurious correlations, we create $K=100$ randomized datasets by randomly swapping word tokens between communities. In principle, randomization breaks the link between individual communities and their contextual word statistics. Consequently, lead-lag relationships retained in the randomized dataset must be attributed to either structural bias or random noise. We filter out such spurious lead-lag relationships by comparing the lead score for a leadership event in the original non-randomized dataset against a set of lead scores for the same event in each of the randomized datasets. Any event in the final set of leadership events for further analysis satisfies the following criterion,
$\textsc{Lead}(e)>\Phi_{.95}\left(\{\textsc{Lead}^{(k)}(e)\}_{k=1}^{K}\right),$ (5)
where the function $\Phi_{.95}(S)$ selects the 95th percentile value of the set $S$. The final output of the entire procedure is a list of events that can be aggregated to produce a weighted directed network between the communities, with the weight on any edge indicating the number of linguistic changes for which the pair of communities are in a lead-lag relationship of statistical significance.
Results
Table 2: Semantic changes Examples of semantic changes with the tweets in which they appear, paraphrased from the original.
Word Span Earlier Tweet Later Tweet
[…]
Our approach offers insight into the changing meaning of conceptual keywords (?) related to political activism and the BLM movement, as well as about the specific communities that helped to spread those changed meanings. Here we discuss several sets of semantic changes identified by our model and the significance of the overall network we construct. Taken together, these results affirm the validity of our model and its underlying community clustering approach, as well as the crucial work done by BLM activists to shape conversations about police violence, antiblack racism, and possible solutions to both. They also affirm the importance of Black celebrities in communicating the ideas of BLM to the general public, as well as the presence of conservative critics who push back against activist discourse–and, at times, co-opt its language. Below we discuss the findings from each phase of our methodological pipeline, building towards a conclusion about the decade-long lead-up to the current political climate in which the words of antiracism themselves have become the subject of legislative attention.
Semantic Changes
The first part of our model identifies the words that demonstrated statistically-significant semantic changes (Table 2). Beginning with general terms related to activism, we find that the term activists undergoes a shift in meaning between 2014-2015, following the murders of Freddie Gray and Sandra Bland, when it has no particular connection to the work of BLM–its near neighbors are terms like “detained,” “targeted,” and “scuffle”–to the summer of 2020, by which point the term is very clearly associated with BLM activists in particular, as indicated by near neighbors including “blackled,” “movements,” “bipoc,” and “advocates.” The term action, similarly, enters the dataset with no specific connotations of activism (its near neighbors are related to TV news channels) but, by the summer of 2020, has acquired specific valences of political change, as evidenced by the near neighbors of “meaningful,” “tangible,” “steps,” and “stand.” In a similar manner, the word active also acquires connotations specific to allyship, as indicated by near neighbors of “learning,” “listening,” and “unlearning,” among others. (The terms allyship and learning undergo similar changes, acquiring each other as near neighbors along similar timeframes).
Terms like “listening” and “learning” point to more specific ways that BLM activists shaped the broader conversation about how to engage in activism around racial justice as well as how to contribute as an ally. But additional terms index specific issues and debates. For example, the term abolish begins with associations with the death penalty and, in 2020, acquires an expanded meaning that encompasses the abolition of the police–a key component of the BLM platform. The term abolishing, similarly, acquires the near neighbors of “defunding,” “reallocating,” “reforming,” and “demilitarization” as the summer of 2020 approaches. We also find theory, as in “critical race theory,” coalescing as a keyword during this time, acquiring associations with both “crt” and “ideology.” Taken together, these word changes point to a clear coalescing around the discourse of racial justice that we have come to associate with the movement.
Figure 5: Linguistic leadership The linguistic leadership of key communities including (a) activists, (b) progressives, (c) Black celebrities, and (d) conservatives with their contributions highlighted. The height of each bar is proportional to the overall number of words a community leads (left) or follows (right). The height of the connecting band is proportional to the number of words the community on the left leads the one on right.
Leadership Network
Unexpected insights begin to emerge when looking at the leadership network overall. To be sure, we observe the influence of the BLM activists (Figure  5(a)), who claim an outsized number of word changes in comparison to most other communities. The progressive community, which sits slightly closer to the center than the activist community, can count a significant number of word changes as well (Figure  5(b)), and it would make sense that the pattern of the progressives would track that of the BLM activists but less strongly. What is perhaps unexpected is that the conservative community-moreso than the center/left news media-is the most substantial follower of both the activist and progressive communities. While we might expect a more gradual path from the left to the center, and then to the right, our method shows how the conservative community engaged directly with the activist and progressive communities from the start.
Another unexpected but not surprising result was the strong role of the community of Black celebrities in communicating the ideas of the BLM movement to the general public (Figure 5(c)). Indeed, Black celebrities claim 24.73% of all of word changes, and are followed by activists, then center/left news media, and conservatives in high proportion. This finding suggests that the direct association with the BLM movement may mark its activists for pushback-and, at times, outright aggression-by the conservative community, the celebrity status of those in the Black celebrity community may protect them from some of the conservative vitriol, and as a result, allow them to communicate the ideas of the activists to a broader public. We see this in their leadership of terms like activism, which transforms from the specific context of social movements to a word that itself indexes a role of social importance, as indicated by the near neighbors of “importance,” “career,” “societal,” and “impact.” We also see this in the terms that one would seem to associate with the BLM movement itself, such as listen. The connotations of learning and unlearning, as well as amplifying the voices of those with direct experience-as indicated by the near neighbors of “learn,” “unlearn,” “amplify,” and “explain,” are communicated to the center/left news media commmunity from the Black celebrities. In addition, we find this community influencing the community of progressives, passing along the changed meaning of theory, discussed above, as well as the more basic concept of equity. Interestingly, the conservative community at times learns directly from the Black celebrity community. In particular, the term nonracist is introduced by this community, and is adopted directly from the conservative community from there.
A final observation has to do with the role of the conservative community in introducing new word meanings-and not just adopting them (Figure 5(d)). While few of these words are conceptual keywords, their overall number is substantial. This finding confirms the consistent and persistent engagement of these ideas by the conservative community-not only in the wake of the summer of 2020, but from the movement’s start. That these words are not conceptual keywords at once confirms prior findings about how conservatives-as well as white people as a group-demonstrate a lead-lag relationship with BLM activists (?) and shows how they engaged at the level of discourse: through the wide array of words that make up a language.
Implications and Limitations
The wide range of statistically significant words in our dataset that are associated with political activism and with the BLM movement in particular, as well as the communities that they were introduced and/or adopted by, confirm the effectiveness of our approach for detecting semantic change and measuring leadership within a large and heterogeneous Twitter community. It also provides an additional layer of evidence for the valuable work of BLM activists, as well as Black celebrities, in advancing the cause of racial justice over the past decade. More recently, as Twitter has become X, and has prompted the scattering of the diverse communities who previously conversed together, we may need to consider additional adaptations to our method of detecting semantic change, just as we will need to reimagine the ways in which activists can spread their messages to a broader public and enlist others in their cause.
As far as the prominence of the conservative community in our study, we see significant implications for our understanding of the emergence of the “anti-woke” and “anti-CRT” agenda that has entered legislation in nine US states and counting (?). Whereas the common narrative is that this movement emerged in the aftermath of the summer of 2020, and has since been fueled by the Republican party, our findings show that there has always been direct conservative engagement with the discourse of Black Lives Matter. Moving forward, this finding may reshape the narrative of “conservative backlash” and, instead, point to a story of continual pushback and even outright aggression that met the movement from its start.
Finally, we would like to highlight a key absence in this project, which is the missing data that might otherwise document the crucial role of the young Black activists who fueled the movement from its start. These regular young people were essential in the movement’s early phases, as they were in organizing the on-the-ground protests that took place in every city where a police killing took place. Qualitative researchers have documented their crucial role (??), as have news reports (?), yet their status as regular young people has meant that they lack the protections granted to celebrities or even to adults. As activists, they have been subjected to harassment and intimidation, and in some cases they have even been killed by the police themselves (??). As early as mid-2015, when Freelon et al released their initial findings on the network structure of #BLM, they found that the young Black activists who had been present in their data in 2014 had disappeared eighteen months later (?). Using the usernames of their young Black activist community as seeds, we found similarly that these users had largely disappeared-save for the few who themselves became famous, and joined the Black celebrity community instead.
In her influential theorization of missing data sets, the artist and educator Mimi Onuoha explains how the gaps in datasets provide “cultural and colloquial hints of what is deemed important,” and further observes how “this lack of data typically correlates with issues affecting those who are most vulnerable in that context” (?). In our #BlackLivesMatter dataset and the findings that it has enabled, these young Black activists constitute this precise form of missing data. Thus as we celebrate our findings for how they affirm the role of the Black Lives Matter movement in shaping a larger public conversation, we must also hold space for what our study cannot confirm, which is the equally crucial role played by the activists whose words are missing from our dataset. Even if their words are unrecorded, they are not unremembered.
Conclusion
In this project we describe a method of modeling semantic leadership across a set of communities associated with the #BlackLivesMatter movement, which has been informed by qualitative research on the structure of social media and of Black Twitter in particular. We describe our bespoke approaches to time-binning, community clustering, and connecting communities over time, as well as our adaptation of state-of-the-art approaches to semantic change detection and semantic leadership induction. We find substantial evidence of the leadership role of BLM activists and progressives, as well as Black celebrities. We also find evidence of the sustained engagement of the conservative community with this discourse, suggesting an alternative explanation for how we arrived at the present moment, in which ”anti-woke” and “anti-CRT” bills are being enacted nation-wide. Finally, we recognize the role of the young Black activists who are missing from our data, and hold space for their contributions even as they cannot be identified through quantitative means.
Appendix
[Uncaptioned image] Figure 6: Top Community Members The most central users of each subcommunity cluster, with color indicating final community grouping. Users with fewer than 3000 followers or with no longer active X accounts are not shown. Note that ranking by in degree leads to certain users (e.g. @realdonaldtrump) being included in communities in which they engage adversarially.
Maxima Additional Candidate Maxima Sep 02, 2014 n/a Dec 25, 2014 Dec 18, 2014 Aug 31, 2015 Mar 14, 2015; Apr 10, 2015; May 04, 2015 Jul 08, 2016 n/a May 27, 2020 n/a Aug 07, 2020 Sep 05, 2020; Sep 25, 2020
Table 3: Maxima Additional candidate maxima that were merged into the final time bin segments.
Community Short Label Black Twitter btwitter Allies/Academics allies BLM Activists activists Progressives progressives Conservatives conserv Black Celebrities bcelebs Mixed Celebrities mcelebs Center/Left Politicians centleftpols Center/Left News Media centleftnews Local News localnews International Users intl
Table 4: Labeling Scheme Table translating between the long and short labels used throughout this paper.
Figure 7: Viable thresholds A knee plot showing the number of matched communities at Jaccard similarity thresholds between 0 and .3 in increments of .0125.